let Z1 =a+bι ,and Z2 =c+ιd then find Z1.Z2
Answer by ikleyn(52855)

.

Z1*Z2 = (a + bi)*(c + di) = ac + bci + adi + bi*di = ac + bci + adi + bd*i^2 = 

      = ac + (bc + ad)i - bd = (ac - bd) + (bc + ad)i.    ANSWER


On the way, I used  i^2 = -1.


Real      part of  Z1*Z2  is  ac - bd.

Imaginary part of  Z1*Z2  is  bc + ad.

Solved.